# The Larger the Better? Improved LLM Code-Generation via Budget Reallocation

### v1

@@ Abstract
It is a common belief that large language models (LLMs) are better than smaller-sized ones. However, larger models also require significantly more time and compute during inference. This begs the question: what happens when both models operate under the same budget? (e.g., compute, run-time). To address this question, we analyze code generation LLMs of various sizes and make comparisons such as running a B model once vs. generating five outputs from a B model and selecting one. Our findings reveal that, in a standard unit-test setup, the repeated use of smaller models can yield consistent improvements, with gains of up to % across five tasks. On the other hand, in scenarios where unit-tests are unavailable, a ranking-based selection of candidates from the smaller model falls short of the performance of a single output from larger ones. Our results highlight the potential of using smaller models instead of larger ones, and the importance of studying approaches for ranking LLM outputs.
@@ 1 Introduction
The common wisdom in deep learning, and language modeling in particular, is that investing more compute leads to improved performance (Kaplan et al., 2020). The standard way of implementing this principle is training larger models. A simpler, yet often overlooked way to increase compute budget is to run a smaller model multiple times, and select the best output using some metric (Chen et al., 2021). In this work we systematically compare between these two approaches: we ask whether, given a fixed compute budget, it is best to run a large model once, or a smaller model multiple times (Figure 1). Our results show that, perhaps surprisingly, given the same compute budget, running 7B or 13B models can not only match the performance of a 70B model, but also substantially surpass it.
@@
Addressing our research question requires a method for selecting the best LLM output from a given set of candidates. In this work we focus on execution based code-generation tasks, which assume the availability of unit-tests (Chen et al., 2021; Austin et al., 2021; Hendrycks et al., 2021). We adapt the widely-used pass@ metric (Kulal et al., 2019), which evaluates a model’s performance on code generation problems by generating outputs and assigning a point if any of them passes all tests. To adapt this metric for our purposes, we consider models of different sizes, and for each generate as many outputs as possible given a fixed compute budget, e.g., floating point operations (FLOPs) or wall-time.
@@
We apply this setup to evaluate the Code Llama (Roziere et al., 2023) model family (B, B, B, and B) across five tasks: HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and the three splits of APPS (Hendrycks et al., 2021). Surprisingly, we find that for the two popular tasks, HumanEval and MBPP, the smaller models (B and B) outperform the larger ones (B and B) by a margin of up to %. Importantly, this is observed using both budget types (FLOPs and wall-time) and across all computation budgets. When considering the challenging APPS benchmark, we find that the B model performs best across almost all budgets, with a consistent margin of % when considering the hardest split—competition.
@@
We then proceed to examine the scenario where unit-tests are unavailable, such as in an IDE code-completion setup. In such cases, an efficient policy is required to select a single solution from all generated ones. We consider a simple LLM ranking policy, which ranks solutions based on the negative log likelihood of the LLM. We experiment with the B model, and rank its outputs using each of the models. Our results show that, as expected, our ranking-based selection improves with the increase in compute budget, and with the size of the ranking LLM. Nonetheless, this procedure still falls short of the performance achieved by running the larger model independently with the same budget.
@@
Our results highlight the potential of using smaller models instead of larger ones, a practice that has many benefits. First, small models are far computationally cheaper to pre-train.111E.g., Llama- B was faster to pre-train compared to the B variant (Touvron et al., 2023). Further, at inference time, they are considerably more hardware-friendly: a B model can be accommodated on a single A GPU, a feat unachievable for a B model (Dettmers et al., 2022). Finally, as we have shown, when controlling for the compute budget, smaller models may actually outperform larger ones.
@@
Our findings also emphasize the importance of developing effective ranking approaches for LLM outputs. This is especially important in cases where no unit-tests or other verification methods are available (Zou et al., 2021; Uesato et al., 2022; Sun et al., 2023). To support such research direction, we release Code Llama B outputs for each example in HumanEval and MBPP—a total of more than M outputs.
@@ 2.1 Standard Code Generation Evaluation
2.2 Comparing LLMs of Different Sizes with a Fixed Budget
Our goal is to compare between LLMs of different sizes, which all have access to the same compute budget. To do so, we allow smaller models, which consume fewer resources, to generate more outputs. As a result, the generation process of both models requires roughly the same amount of compute.
@@ 3.2 Metrics
Computing the pass@ and pass@ metrics requires an estimation of the flops() and time() functions from Equations 2 and 3. To estimate the FLOPs usage of each model we use the calflops library (xiaoju ye, 2023), with input sequence length of . For the estimation of wall-time usage of a model, we measure wall-time while assuming optimal throughput utilization of the hardware. Specifically, we use a node of A GPUs, optimize the batch size per model and measure the time it takes each model to generate a subset of k examples from our datasets. We report the results in Table 1, for readability we also report the normalized factor with respect to the B model.
@@ 3.3 Experiments
We use the Code Llama family (Roziere et al., 2023), a finetuned version of Llama (Touvron et al., 2023) for code generation. Code Llama comes in various sizes, which we use for our experiments: B, B, B and B.
We follow Roziere et al. (2023), and use a zero-shot setting for HumanEval, a -shot prompting strategy for MBPP and -shot prompts for APPS, and limit the generation length to // tokens for HumanEval/MBPP/APPS. For the sampling process, we use nucleus sampling (Holtzman et al., 2019) with top- and a temperature of for HumanEval/MBPP/APPS, with all models sizes (Roziere et al., 2023). Additionally, we provide the pass@1 results using a greedy decoding method for all models.
To compare models in varying sizes, we select the maximal number of generations for each model with respect to the values in Table 1. Specifically, for the smaller benchmarks, HumanEval and MBPP, we generate answers for the B/B/B/B models, respectively. For the larger benchmarks, the three splits of APPS, we use . To get a robust estimation of these measures, we follow Chen et al. (2021) and Roziere et al. (2023), and report for all benchmarks a maximal value of for the pass@ metric.
@@ 4 Small Models Outperform Large Ones under a Fixed Compute Budget
Results for HumanEval and MBPP are presented in Figures 2 and 3, respectively. We first note that, as expected, the pass@ metric improves both with model scale, and with the number of generations (Figures 2(a) and 3(a)). However, perhaps surprisingly, when considering the pass@ and pass@ metrics (Figures 2(b), 2(c), 3(b) and 3(c)), we see a different trend—given a fixed compute budget, smaller models yield better results than larger ones. Specifically, the B/B models outperform the larger models across all compute budgets. Particularly, in the small budget regime (up to normalized FLOPs units and wall-time units) the performance gap widens to —.
Another way of looking at our results is by observing that smaller models match the performance of larger ones using substantially lower budgets. For instance, in HumanEval, the B and B models achieve a score of using one quarter of the time it takes the larger models to reach that score. Finally, we compare small models to greedy decoding with larger models, which generally performs better than sampling. We observe that even in this setup, using the smaller models several times is equivalent or preferable in all cases.
We next turn to discuss the results over the three splits of the APPS benchmark (Figures 4, 5 and 6). We first consider the B model, and observe the same trends as in HumanEval and MBPP: this model achieves the best performance in almost all fixed compute budgets. Specifically for the competition split (Figures 6(b) and 6(c)), the most challenging APPS split, the B model outperforms all other models in all compute budgets, with a consistent margin of from the B model when considering the wall-time budget. We further observe that the B model achieves similar or better performance as the greedy approach of all models in all three splits. Finally, when fixing the performance, we observe that the B model is – times more efficient than the B model (both for FLOPs and wall-time).
@@
Another intriguing aspect of our research question is the performance of LLMs under a fixed budget when unit-tests are unavailable, and a single selection among several generations must be made. We delve into this topic in the following section.
@@ 5 Evaluating Code Generation without Unit-tests
We examine the scenario where unit-tests are not available (e.g., IDE code-completion setup). In this case, an efficient selection policy strategy may be used to select one answer from the model’s generations. In the previous cases (Section 2), unit-tests served as this policy. Here we investigate using ranking as a selection policy. In Section 5.1 we show how to estimate the performance of a model given a ranking strategy, and in Section 5.2 we analyze the performance of larger models as rankers for a small model.
@@ 5.2 Large Language Models as Rankers
We examine the usage of LLMs as rankers. To produce a ranking order over a set of generations, we use the averaged Negative Log Likelihood (NLL) the LLM assigns to each generation (excluding the prompt), and rank the generations according to that score. It should be noted that extracting the NLL of a model over a given generation can be done in a parallel manner (i.e., non-autoregressively), which is substantially more efficient compared to traditional token-by-token generation. The score given by a model to a generation given a prompt is:
@@
Our results are presented in Figure 8. As can be seen, using LLMs as rankers over generations obtained from smaller models improves performance. Interestingly, we observe that using a B model as a ranker for itself can enhance its generation even further than the greedy approach, albeit with the cost generating several outputs. We also find that using larger models as rankers results in better perfomance.
When considering a fixed compute budget, we find that it is sometimes comparable to use LLMs as rankers instead of sampling from them, as can be seen with the B and B models. However, this is not the case for the greedy approach which consistently outperforms ranking multiple generations from a smaller model given a fixed compute budget. In summary, there remains a gap to bridge between using LLMs as rankers for smaller models and using them as generators. To further promote this line of research, we release the generations per example produced by the B model for both HumanEval and MBPP (a total of generations).
@@ 6 Related Work
was found to be one of the key elements in the success of LLMs (Dehghani et al., 2023; Gu et al., 2023; Hassid et al., 2024; Rae et al., 2021; Chowdhery et al., 2023; Touvron et al., 2023), with Wei et al. (2022) demonstrating how specific abilities emerge mainly after reaching a specific scale. The way language models behave when they are scaled up and their ability to adjust have been a significant factor in the creation of LLMs (Hernandez et al., 2021). Kaplan et al. (2020) investigated the optimal model size to train for a given compute budget, while Hoffmann et al. (2022) demonstrated how scaling both model and dataset sizes improves performance across various tasks. Clark et al. (2022) analyzed the scaling properties of mixture-of-experts models, showing that scaling with the number of experts diminishes as model size increases. Recently, Gadre et al. (2024) provided a scaling laws analysis considering downstream tasks rather than next-token prediction loss. They related the perplexity of a language model to its downstream task performance via a power law and used it to predict the top-1 error averaged over the evaluated downstream tasks. Our work differs from all of the above, as we do not claim to provide new scaling laws but rather suggest that when fixing the budget, smaller models can provide comparable or superior results to larger ones. Finally, a very recent work by Shi et al. (2024) showed that small vision models can outperform larger models if they are fed with multiple scales of the image. Our approach, which generates multiple text outputs from a small model, is similar in spirit to theirs.
LLM verifiers and rankers
@@
is a growing trend, which leverages LLMs to verify and rank generations obtained from weaker and smaller models (Cobbe et al., 2021b; Uesato et al., 2022; Saha et al., 2024; Havrilla et al., 2024). Both Cobbe et al. (2021b) and Uesato et al. (2022) leveraged an external classifier to rank LLM outputs. Specifically in both setups the authors proposed to generate many candidate solutions and select the one ranked highest by the verifier. The authors demonstrated the applicability of using such verifiers in solving math word problems (Cobbe et al., 2021a). Qin et al. (2023) demonstrated that LLMs can serve as efficient text rankers when considering pairwise ranking. Another line of work leveraged LLMs to evaluate the quality of smaller models (Saha et al., 2024; Dubois et al., 2023; Zheng et al., 2023; Oren et al., 2024). Although providing promising alternative, such evaluation suffers from biases in the larger model (Zheng et al., 2023) and reliance on hand-designed evaluation plans that impact the method’s ability to generalize (Liu et al., 2023). Large models also serve as verifiers of small ones in a speculative decoding setup, with the goal of speeding-up LLM generation (Leviathan et al., 2023; Kim et al., 2023; Chen et al., 2023). In this work we explore the potential of LLMs as selectors of the best output of a smaller model in a fixed budget setup. Similarly to ours, Li et al. (2024) found that smaller sized LMs (B parameters) already exhibit strong mathematical abilities when selecting the best response from different generations. When considering code generation models, AlphaCode Team (2023) presented impressive results on challenging coding contests tasks while generating M samples, and later on filtering and ranking them using Gemini-Pro LLM (Team et al., 2023). Lastly, Dou et al. (2024) proposed a method to improve code-generation models by learning a policy model using reinforcement learning methods. This is done by sampling several solutions from the LLM and scoring them using a feedback obtained from the compiler.

### v2

@@ Abstract
It is a common belief that large language models (LLMs) are better than smaller-sized ones. However, larger models also require significantly more time and compute during inference. This begs the question: what happens when both models operate under the same budget? (e.g., compute, run-time). To address this question, we analyze code generation LLMs of various sizes and make comparisons such as running a B model once vs. generating five outputs from a B model. We consider a standard unit-test setup, which can be used to select the correct output from the smaller model. Our findings reveal that the repeated use of smaller models can yield consistent improvements, with gains of up to % across five tasks. On the other hand, in scenarios where unit-tests are unavailable, a ranking-based selection of candidates from the smaller model falls short of the performance of a single output from larger ones. Our results highlight the potential of using smaller models instead of larger ones, and the importance of studying approaches for ranking LLM outputs.111 Data is avalible at https://github.com/slp-rl/budget-realloc
@@ 1 Introduction
A common wisdom in deep learning, and language modeling in particular, is that investing more compute leads to improved performance (Kaplan et al., 2020). The standard way of implementing this principle is training larger models. A simpler, yet often overlooked way to increase compute budget is to run a smaller model multiple times, and select the best output using some metric (Chen et al., 2021). In this work we systematically compare between these two approaches: we ask whether, given a fixed compute budget, it is best to run a large model once, or a smaller model multiple times (Figure 1). Our results show that, perhaps surprisingly, given the same compute budget, running 7B or 13B models can not only match the performance of a 70B model, but also substantially surpass it.
@@
Addressing our research question requires a method for selecting the best LLM output from a given set of candidates. In this work we focus on execution based code-generation tasks, which assume the availability of unit-tests (Chen et al., 2021; Austin et al., 2021; Hendrycks et al., 2021). We consider the widely-used pass@ metric (Kulal et al., 2019), which evaluates a model’s performance on code generation problems by generating outputs and assigning a point if any of them passes all tests. To adapt this metric for our purposes, we take models of different sizes, and for each generate as many outputs as possible given a fixed compute budget, e.g., floating point operations (FLOPs) or wall-time.
@@
We apply this setup to evaluate the Code Llama (Roziere et al., 2023) model family (B, B, B, and B) across five tasks: HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and the three splits of APPS (Hendrycks et al., 2021). For the HumanEval and MBPP benchmarks, we additionally use the recent Llama-3 (AI@Meta, 2024) model family (B and B). Surprisingly, we find that for the two popular tasks, HumanEval and MBPP, the smaller models (B, B and B) outperform the larger ones (B and B) by a margin of up to %. Importantly, this is observed using both budget types (FLOPs and wall-time) and across all computation budgets. When considering the challenging APPS benchmark, we find that the B model performs best across almost all budgets, with a consistent margin of % when considering the hardest split—competition.
@@
We then proceed to examine the scenario where unit-tests are unavailable, such as in an IDE code-completion setup. In such cases, an efficient policy is required to select a single solution from all generated ones. We consider a simple LLM-based policy, which ranks solutions based on the negative log likelihood of the LLM. We also augment this policy with a variant of a recent ranking approach—LEVER (Ni et al., 2023). We experiment with the B model, and rank its outputs using each of the models. Our results show that, as expected, ranking-based selection improves with the increase in compute budget, and with the size of the ranking LLM. Nonetheless, this procedure still falls short of the performance achieved by running the larger model independently with the same budget.
@@
Our results highlight the potential of using smaller models instead of larger ones, a practice that has many benefits. First, small models are far computationally cheaper to pre-train.222E.g., Llama- B was faster to pre-train compared to the B variant (Touvron et al., 2023). Further, at inference time, they are considerably more hardware-friendly: a B model can be accommodated on a single A GPU, a feat unachievable for a B model (Dettmers et al., 2022). Finally, as we have shown, when controlling for the compute budget, smaller models may actually outperform larger ones.
@@
Our findings also emphasize the importance of developing effective ranking approaches for LLM outputs. This is especially important in cases where no unit-tests or other verification methods are available (Zou et al., 2021; Uesato et al., 2022; Sun et al., 2023). To support such research direction, we release Code Llama B outputs for each example in HumanEval and MBPP—a total of more than M outputs.11footnotemark: 1
@@ 2.1 Standard Code Generation Evaluation
2.2 Comparing LLMs of Different Sizes under a Fixed Budget
Our goal is to compare between LLMs of different sizes under a fixed compute budget. To do so, we allow smaller models, which consume fewer resources, to generate more outputs. This results in models of different sizes requiring roughly the same amount of compute.
@@ 3.2 Metrics
Computing the pass@ and pass@ metrics requires an estimation of the flops() and time() functions from Equations 2 and 3. To estimate FLOPs usage, we use the calflops library (xiaoju ye, 2023), with input sequence length of . We measure wall-time while assuming optimal throughput utilization of the hardware. Specifically, we use a node of A GPUs, optimize the batch size per model and measure the time it takes each model to generate a subset of 1k examples from our datasets. We report the Code Llama results in Table 1, for readability we also report the normalized factor with respect to the B model.333Llama-3 B/B presents similar usage to Code Llama B/B, with a difference of up to 7%.
@@ 3.3 Experiments
We experiment with the Code Llama family (Roziere et al., 2023), a finetuned version of Llama (Touvron et al., 2023). Code Llama comes in various sizes, which we use for our experiments: B, B, B and B. For the smaller benchmarks, HumanEval and MBPP, we also consider the Llama-3 family (B and B).
We follow Roziere et al. (2023), and use a zero-shot setting for HumanEval, a -shot prompting strategy for MBPP and -shot prompts for APPS, and limit the generation length to // tokens for HumanEval/MBPP/APPS. For the sampling process, we use nucleus sampling (Holtzman et al., 2019) with top- and a temperature of for HumanEval/MBPP/APPS, with all models sizes (Roziere et al., 2023). Finally, we also report, the pass@1 results using a greedy decoding method for all models.
To compare models in varying sizes, we select the maximal number of generations for each model with respect to the values in Table 1. Specifically, for the smaller benchmarks, HumanEval and MBPP, we generate answers for the -B/B/B/B models, respectively. For the larger benchmarks, the three splits of APPS, we use . To get a robust estimation of these measures, we follow Chen et al. (2021) and Roziere et al. (2023), and report for all benchmarks a maximal value of for the pass@ metric, while using all available unit-tests.
@@ 4 Small Models Outperform Large Ones under a Fixed Compute Budget
Results for HumanEval and MBPP using the Code Llama models are presented in Figures 2 and 3, respectively.444Tables 2 and 3 in Appendix B presents detailed results. The corresponding results for the Llama-3 models can be found in Figures 10 and 11 (Appendix A). We first note that, as expected, the pass@ metric improves both with model scale, and with the number of generations (sub-figure (a) in all figures). However, perhaps surprisingly, when considering the pass@ and pass@ metrics (sub-figures (b) and (c)), we see a different trend—given a fixed compute budget, smaller models yield better results than larger ones. Specifically, the B/B/B models outperform the larger models across all compute budgets. Particularly, in the small budget regime (up to normalized FLOPs units and wall-time units) the performance gap reaches —.
Another way of looking at our results is by observing that smaller models match the performance of larger ones using substantially lower budgets. For instance, in HumanEval, the Code Llama B and B models achieve a score of using one quarter of the time it takes the larger models to reach that score. This efficiency gap further increases with the Llama-3 models (Figure 10(c)). Finally, we compare small models to greedy decoding with larger models, which generally performs better than sampling. We observe that even in this setup, using the smaller models several times is equivalent or preferable in all cases.
We next turn to discuss the Code Llama results over the three splits of the APPS benchmark (Figures 4, 5 and 6).555Table 4 in Appendix B presents detailed results. We first consider the B model, and observe the same trends as in HumanEval and MBPP: this model achieves the best performance in almost all fixed compute budgets. Specifically for the competition split (Figures 6(b) and 6(c)), the most challenging APPS split, the B model outperforms all other models in all compute budgets, with a consistent margin of from the B model when considering the wall-time budget. We further observe that the B model achieves similar or better performance as the greedy approach of all models in all three splits. Finally, when fixing the performance, the B model is – times more efficient than the B model (both for FLOPs and wall-time).
@@
Our results indicate that small models can match or even outperform large ones under a fixed compute budget, assuming the availability of unit-tests. An intriguing aspect of our research question is what happens when unit-tests are unavailable, and a single selection among several generations must be made. We delve into this topic in the following section.
@@ 5 Evaluating Code Generation without Unit-tests
We examine the scenario where unit-tests are not available (e.g., IDE code-completion setup). In this case, an efficient selection policy strategy may be used to select one answer from the model’s generations. In the previous cases (Section 2), unit-tests served as this policy. Here we investigate using ranking as a selection policy. In Section 5.1 we show how to estimate the performance of a model given such a strategy, and in Section 5.2 we analyze the performance of larger models as rankers for a small model.
@@ 5.2 Large Language Models as Rankers
We examine the usage of LLMs as rankers. To produce a ranking order over a set of generations, we use the averaged Negative Log Likelihood (NLL) the LLM assigns to each generation (excluding the prompt), and rank the generations according to that score. It should be noted that extracting the NLL of a model over a given generation can be done in a parallel manner (i.e., non-autoregressively), which is substantially faster than traditional token-by-token generation. The score given by a model to a generation given a prompt is:
@@
Our results are presented in Figure 8. As can be seen, using LLMs as rankers over generations obtained from smaller models improves performance. Interestingly, we observe that using a B model as a ranker for itself can enhance its generation even further than the greedy approach, albeit with the cost generating several outputs. We also find that using larger models as rankers results in better perfomance. When considering a fixed compute budget, we find that it is sometimes comparable to use LLMs as rankers instead of sampling from them, as can be seen with the B and B models. However, this is not the case for the greedy approach which consistently outperforms ranking multiple generations from a smaller model given a fixed compute budget.
To further check the use of external verifiers, we integrate the LEVER verifier model (Ni et al., 2023) with the Code Llama models. The LEVER approach aims to enhance code generation by learning to verify generated programs. The full LEVER pipeline involves using the NLL produced by the code generation model, error pruning based on execution, and a verifier trained on code generations with execution results. However, since we assume that no tests are available in our setting, execution pruning and execution results cannot be used. LEVER released a trained verifier over the MBPP benchmark, which we use along with the NLL scores of each model. As shown in Figure 9, the LEVER verifier does not improve the results in the test-less setting, which is expected given that one of the main components of the approach relies on execution over unit-tests.
In summary, there remains a gap to bridge between using LLMs as rankers for smaller models and using them as generators. To further promote this line of research, we release the generations per example produced by the B model for both HumanEval and MBPP (a total of generations).
@@ 6.1 Model Scaling
Model scaling was found to be one of the key elements in the success of LLMs (Dehghani et al., 2023; Gu et al., 2023; Hassid et al., 2023; Rae et al., 2021; Chowdhery et al., 2023; Touvron et al., 2023), with Wei et al. (2022) demonstrating how specific abilities emerge mainly after reaching a specific scale. The way language models behave when they are scaled up and their ability to adjust have been a significant factor in the creation of LLMs (Hernandez et al., 2021). Kaplan et al. (2020) investigated the optimal model size to train for a given compute budget, while Hoffmann et al. (2022) demonstrated how scaling both model and dataset sizes improves performance across various tasks. Clark et al. (2022) analyzed the scaling properties of mixture-of-experts models, showing that scaling with the number of experts diminishes as model size increases. Recently, Gadre et al. (2024) provided a scaling law analysis considering downstream tasks rather than next-token prediction loss. They related the perplexity of a language model to its downstream task performance via a power law and used it to predict the top-1 error averaged over the evaluated downstream tasks. Our work differs from all of the above, as we do not claim to provide new scaling laws but rather suggest that when fixing the budget, smaller models can provide comparable or superior results to larger ones.
Recent studies by Shi et al. (2024) and Mei et al. (2024) have demonstrated that under constrained compute budgets, smaller vision models can surpass their larger counterparts. Specifically, Shi et al. (2024) found advantages in using multiple image scales, whereas Mei et al. (2024) observed that smaller diffusion models perform better than larger ones when the compute budget is fixed. Our approach, which generates multiple text outputs from a small model, aligns with these findings.
6.2 Verifiers and Rankers
LLM verifiers and rankers is a growing trend, which leverages LLMs to verify and rank generations obtained from weaker and smaller models (Cobbe et al., 2021b; Uesato et al., 2022; Saha et al., 2024; Havrilla et al., 2024). Both Cobbe et al. (2021b) and Uesato et al. (2022) leveraged an external classifier to rank LLM outputs. Specifically, in both setups the authors proposed to generate many candidate solutions and select the one ranked highest by the verifier. The authors demonstrated the applicability of using such verifiers in solving math word problems (Cobbe et al., 2021a). Qin et al. (2023) demonstrated that LLMs can serve as efficient text rankers when considering pairwise ranking.
Another line of work leveraged LLMs to evaluate the quality of smaller models (Saha et al., 2024; Dubois et al., 2023; Zheng et al., 2023; Oren et al., 2024). Although providing a promising alternative, such evaluation suffers from biases in the larger model (Zheng et al., 2023) and reliance on hand-designed evaluation plans that impact the method’s ability to generalize (Liu et al., 2023). Large models also serve as verifiers of small ones in a speculative decoding setup, with the goal of speeding-up LLM generation (Leviathan et al., 2023; Kim et al., 2023; Chen et al., 2023). It is also common to distill knowledge from a large model into a smaller one in order to improve efficiency (Hinton et al., 2015; Sanh et al., 2019; Xu et al., 2024), see Treviso et al. (2023) for a survey on efficient methods in NLP.
@@ 6.2 Verifiers and Rankers
In this work, we explore the potential of LLMs as selectors of the best output of a smaller model in a fixed budget setup. Similarly to ours, Li et al. (2024) found that smaller sized LMs (B parameters) already exhibit strong mathematical abilities when selecting the best response from different generations. When considering code generation models, AlphaCode Team (2023) presented impressive results on challenging coding contests tasks while generating M samples, and later on filtering and ranking them using Gemini-Pro LLM (Team et al., 2023). Dou et al. (2024) proposed a method to improve code-generation models by learning a policy model using reinforcement learning methods. Lastly, Shi et al. (2022) and Ni et al. (2023) used execution feedback in order to filter code-generations, while Shi et al. (2022) used non-learned approaches, Ni et al. (2023) trained an external verifier on top of the generation and the execution feedback.